OpenStudy (anonymous):

Evaluate the area of the region between the following curves for x in (-1,1) y=4x^2+5 y=-3

OpenStudy (anonymous):

looking to integrate right?

OpenStudy (anonymous):

if that is how you find the area yes!

OpenStudy (anonymous):

that is integrate \[\int_{-1}^1(4x^2+8)dx\] you know how to do this?

OpenStudy (anonymous):

yes and i tried that but it said that answer wasnt right :/

OpenStudy (anonymous):

what did you get as your answer?

OpenStudy (anonymous):

18.66666667

OpenStudy (anonymous):

ah that is why

OpenStudy (anonymous):

not sure how you got it, but lets do it the easy way \[\int_{-1}^14x^2dx+\int_{-1}^18dx\] second part is \(2\times 8=16\) from your eyeballs.

OpenStudy (anonymous):

first part need the anti derivative of \(4x^2\) which is \(\frac{4}{3}x^3\)

OpenStudy (anonymous):

soo how should i calculate the answer?

OpenStudy (anonymous):

oh damn i get the same thing!!!

OpenStudy (anonymous):

sorry i thought something was different is this an on line homework?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer is \(\frac{56}{3}\) if it is on line maybe you should just try the fraction

OpenStudy (anonymous):

yes thats right thank you!!!!

OpenStudy (anonymous):

what system are you using? it is possible that it doesn't like 18.33333...

OpenStudy (anonymous):

you are welcome, but you did all the work. i just helped with the input

OpenStudy (anonymous):

would you know this one? y=e^x+5 y=9 ??

OpenStudy (anonymous):

given no constraints

OpenStudy (anonymous):

what exactly is the question?

OpenStudy (anonymous):

area between the curves?

OpenStudy (anonymous):

that can't be right because there is no area between the curves what does the question ask?

OpenStudy (anonymous):

yes thats what it is asking for

OpenStudy (anonymous):

are you sure? because it doesn't make sense really. the region is not bounded in any way,

OpenStudy (anonymous):

\(y=9\) is just a horizontal line, and \(e^x+5\) is an exponential curve looks like this http://www.wolframalpha.com/input/?i=y%3De^x%2B5%2Cy%3D9

OpenStudy (anonymous):

so there is no region between the curves maybe you could post the problem

OpenStudy (anonymous):

i know thats why its given me trouble byt thx for tryin

OpenStudy (anonymous):

yw